OpenStudy (anonymous):

Given that f(x) = x2 + 3x + 6 and g(x) = x-3 over 2, solve for f(g(x)) when x = 1 SOMEBODY PLEASE EXPLAIN HOW TO SOLVE

OpenStudy (dumbcow):

what is g(1) ?

OpenStudy (anonymous):

Given that f(x) = x2 + 3x + 6 and g(x) = the quantity of x plus two, over three, solve for f(g(x)) when x = 1

OpenStudy (anonymous):

thats the question

OpenStudy (dumbcow):

no sorry i understand the question, i am asking you if you can evaluate g(1) the goal is to find f(g(1))

OpenStudy (anonymous):

thats where im confused idk where to start..

OpenStudy (dumbcow):

if i say h(x) = x - 5 , evaluate h(3) can you do that?

OpenStudy (anonymous):

h(3) = 3 - 5 ?

OpenStudy (dumbcow):

yep good , which is -2 h(3) = -2 do same thing with function g(x) at x=1 g(1) = ?

OpenStudy (anonymous):

g(1) = 1-5 ? what is (g)

OpenStudy (dumbcow):

g is name of the function g(x) is defined in the problem, i used the letter "h" because i made up a new function

OpenStudy (dumbcow):

\[g(x) = \frac{x-2}{3}\]

OpenStudy (dumbcow):

** sorry (x-3)/2 whatever you know what it is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

g(1) = functoin 1 ?

OpenStudy (dumbcow):

??

OpenStudy (anonymous):

im so confused

OpenStudy (dumbcow):

why? just plug 1 in for x into g(x) ok so you get some number g(1) then you plug than number into f(x) to get f(g(1)) this new number is the answer

OpenStudy (anonymous):

so g(1) = 1 ?

OpenStudy (dumbcow):

not sure you confused me about what g(x) is lol

OpenStudy (anonymous):

g(x) = 1 right

OpenStudy (dumbcow):

yes if g(x) = (x+2)/3

OpenStudy (anonymous):

Given that f(x) = x2 + 3x + 6 and g(x) = the quantity of x plus two, over three, solve for f(g(x)) when x = 1

OpenStudy (anonymous):

so how do i do this one

OpenStudy (dumbcow):

haha you are doing it ok g(1) = 1 now evaluate f(1)

OpenStudy (anonymous):

1

OpenStudy (anonymous):

f(1) = 1

OpenStudy (dumbcow):

nope f(x) = x^2 +3x +6 f(1) = 1^2 +3(1) +6